OpenStudy (anonymous):

Can someone show me how to graph xy = 15?

OpenStudy (e.mccormick):

Find y, graph a few points, like the y intercept. and watch out for x=0.

OpenStudy (e.mccormick):

Oh, and finding the y intercept is actually a bit of a trick.... you may notice there is not one. There is also not an x intercept.

OpenStudy (e.mccormick):

When you solve for y, you get: \(y=\frac{15}{x}\) Any x that evenly divides 15 is a good point to use. So \(x=\{-3,-1.5,1.5,3\}\) give useful points.

OpenStudy (anonymous):

Okay.

OpenStudy (e.mccormick):

Then play connect the dots. Remember, you can't cross either the x or y axis.

OpenStudy (e.mccormick):

(1,15), (15,1), (-1,-15), (-15,-1) are also there, so you have an idea what happens to the tails.

OpenStudy (e.mccormick):

The lack of an x or y intercept is one of the big clues to the shape. Do you know what that means? What quadrants your points are in is the other part of this clue.

OpenStudy (e.mccormick):

Yep! https://www.desmos.com/calculator/n7xd8g8ojy

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Could you help me with one more?

OpenStudy (e.mccormick):

Sure, but first... one concept from this type graph becomes super important later in math. Can you guess why the graph shoots up and down as it gets close to the origin?

OpenStudy (anonymous):

Not exactly sure.

OpenStudy (anonymous):

Because it is undefined?

OpenStudy (e.mccormick):

\(y=\frac{15}{x}\) Well, what if x is really small? Close to 0. Say \(x=\frac{1}{100}\) \[y=\frac{15}{\frac{1}{100}}\\ y=\frac{15}{1}\cdot\frac{100}{1}\\ y=1500\]

OpenStudy (anonymous):

Oh, okay.

OpenStudy (e.mccormick):

The undefined part then becomes what they call an asymptote. Vertical asymptotes are lines that can not be crossed.

OpenStudy (anonymous):

Alright. One more problem?

OpenStudy (e.mccormick):

Sure.

OpenStudy (anonymous):

Okay. \[y = \frac{ -18 }{ x }\]

OpenStudy (e.mccormick):

Same principal.

OpenStudy (e.mccormick):

It just changes what quads because of the -.

OpenStudy (e.mccormick):

The important thing for easy graphing of these are factors of 18. Half points are not bad, but sometimes when you get down to weird numbers it just gets hard to graph nicely. 18 9 6 3 2 1 are all factors of 18. So they make the "easy" points this time. 1.8 is OK, just not as easy to graph as whole numbers and halfs.

OpenStudy (anonymous):

Okay.

OpenStudy (e.mccormick):

And you will get something like:Not to scale this time, but I am sure that you get the point by now.

OpenStudy (anonymous):

For some reason -1.8,-10 and 1.8,10 aren't plotting right, I know you said it would be a bit harder.

OpenStudy (e.mccormick):

(-1.8,10),(1.8,-10)

OpenStudy (anonymous):

Yeah, I just realized I did it wrong, thanks.

OpenStudy (e.mccormick):

np. Have fun!

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (e.mccormick):

Yah, not too terrible once you have the basic concept.